OpenStudy (anonymous):

If Andy and Billy work together, they can complete a project in 3.6 hours. If each of them works individually, Andy needs 3 hours less than Billy to complete the project. How long does Billy take to complete the project on his own?

OpenStudy (dmezzullo):

\[{\rm Time =}\dfrac{ab}{a + b}\]U have \(a = b - 3\)

OpenStudy (dmezzullo):

\[\dfrac{b(b - 3)}{b + b - 3} = 3.6\]solve?

OpenStudy (anonymous):

b=1.2 or b=9 ?

hartnn (hartnn):

those are correct values of b.

hartnn (hartnn):

which value of 'b' you can eliminate ?

OpenStudy (dmezzullo):

HINT: \(a = b -3\) but one of the \(a\)'s will be negative.

OpenStudy (anonymous):

a=9-3=6 a=1.2-3=-1.8(rej.)

OpenStudy (anonymous):

so i should get the answer a=6 and b=9. Therefore, Billy takes 9 hours to complete the project on his own. Right?

OpenStudy (dmezzullo):

thats right

OpenStudy (dmezzullo):

yes!

OpenStudy (anonymous):

thanks

hartnn (hartnn):

\(\huge \color{green}\checkmark\)